# Kishu: Time-Traveling for Computational Notebooks

### v1

@@ 8. Related Work
Notebook processes on Jupyter-based platforms can be checkpointed by serializing the data in the session state with various libraries (Foundation, 2023f, g, e, d; MongoDB, 2023; MessagePack, 2024; DuVall, 2015; cloudpipe, 2024). There exists a variety of checkpointing tools built on these libraries: On-disk KV-stores can save individual variables (Team, 2023c; shove, 2024; 0xnurl, 2024; Python, 2024; Foundation, 2024b; chest, 2024), DumpSession (Foundation, 2023h) writes dumps containing the serialized bytestring of the session state, ElasticNotebook (Li et al., [n.d.]) combines data storage/loading with cell replay for optimized session migration, and Tensorflow (Google, 2023b) and Pytorch (et al., 2018) offer periodical checkpointing during ML model training. These works do not checkpoint nor restore incrementally, or come with limitations/require significant user effort: Dill and ElasticNotebook always checkpoints the entire state, and their checkpoint files must be loaded in their entirety for restoration. Tensorflow (Google, 2023b) and Pytorch (et al., 2018)’s checkpointing are limited to objects in their respective libraries. While KV stores can store and load individual variables, they do not include delta detection and inter-variable reference preservation, which must be manually handled by the user. In comparison, Kishu addresses all limitations by performing low-overhead incremental checkpointing and restoration while featuring shared reference preservation (i.e., correctness), and works with almost all data science libraries (§ 7.2).
Memory Snapshotting for Checkpoint/Restore
@@
There exists a variety of OS-level checkpointing tools that support incrementally checkpointing a process for later restoration (CRIU, 2023; Ansel et al., 2009; Chen et al., 1997; Plank et al., 1994; Garg et al., 2018; Jain and Cooperman, 2020; Li and Lan, 2010). Typically, these tools create incremental checkpoints by identifying and storing dirty memory pages, then piece back together the process image (across multiple files) for restoration (CRIU, 2023). The main limitations of these tools is (1) large incremental checkpoint file sizes resulting from coarse page-level deltas (Elnozahy et al., 2002), (2) inability to checkpoint across multiple processes (CRIU, 2024), and (3) they can only restore a process from scratch: while we found a patent (Neary et al., 2007) and paper (Ferreira et al., 2011) addressing these limitations enabling OS-level checkpointing for multiprocessing jobs and sub-memory-page granularity incremental checkpointing, respectively, we have been unable to locate working implementations. In comparison, for notebook states, Kishu is able to achieve significantly lower checkpoint overheads via finer Co-variable granularity deltas, checkpoint multiprocessing and off-CPU notebooks via application-level instructions, and fast incremental restore with minimal data loading via state difference detection and leveraging existing data in the process/kernel.
Reverse Debugging
Reverse debugging programs periodically checkpoint the session state, such that a previous intermediate program state can be returned to by loading the appropriate checkpoint and then replaying certain program executions (GDB, 2022; Engblom, 2012; Pothier and Tanter, 2009; Bellard, 2005; VMWare, 2009). Specifically for Python, IPyFlow (Shankar et al., 2022), IncPy (Guo and Engler, 2011), and PyTrace (PyCrunch, 2020) memoize cell and/or function execution results, which can be returned to later for reverse debugging. To the best of our knowledge, existing reverse debuggers do not checkpoint incrementally: users have to balance high overhead from frequent checkpointing and long restore times (via replay) from sparse checkpointing. In comparison, Kishu performs low-overhead incremental checkpointing while enabling fast incremental restore for notebook states.
Lineage Tracing in Computational Notebooks
Lineage tracing in notebook systems has been recently been explored for multi-version notebook replay (Manne et al., 2022), recommending notebook interactions (Macke et al., 2020), creating reproducible notebook containers (Ahmad et al., 2022), and session state migration (Li et al., [n.d.]; Cunha et al., 2021). Tracing methods can be divided into two categories: static code analysis (Foundation, 2023a) or live code instrumentation (GDB, 2022). Tracers in notebook systems need to balance the usage of the two: using static analysis too much will result in many false positives (e.g., due to control flows) (Li et al., [n.d.]), while live analysis can incur prohibitive overhead, especially when tracing at a fine-grained (e.g., object) level (Shankar et al., 2022; Macke et al., 2020). In Kishu, we propose the Co-variable granularity and its accompanying tracing technique using the VarGraph structure: by solely utilizing live analysis at a sufficient granularity (for our use case), we simultaneously minimize false positives while avoiding high overhead from fine-grained live analysis.
@@ 9. Conclusion
In this work, we have proposed Kishu, a new computational notebook system that offers efficient and fault-tolerant time-traveling between arbitrary notebook states. To achieve this, Kishu captures the evolution of session states at the novel Co-variable granularity, enabling time and space-efficient incremental checkpointing of state deltas, which Kishu uses in conjunction with accurate state delta identification to perform incremental checkout with minimal data loading. Its core contributions include (1) low-overhead state delta detection, (2) branch-based session state versioning, and (3) efficient time-traveling with high generalizability—preserving complex inter-variable dependencies and fault tolerance—handling missing or corrupted data through fallback recomputation. We have demonstrated that Kishu is compatible with 146 object classes popular from data science libraries, and can reduce incremental checkpoint storage overhead and checkout time by up to 4.55$\times$ and 9.02$\times$, respectively, on real-world data science notebooks.

### v2

@@ 8.3. Checkpointing Notebook Objects
Data in IPython-based (e.g., Jupyter) notebook session states can be saved with serialization libraries (Foundation, 2023f, g, e, d; MongoDB, 2023; MessagePack, 2024; DuVall, 2015; cloudpipe, 2024), on which a variety of existing checkpointing tools are built: On-disk KV-stores save individual variables (Team, 2023c; shove, 2024; 0xnurl, 2024; Python, 2024; Foundation, 2024b; chest, 2024), DumpSession (Foundation, 2023h) saves the session state in bulk, and ElasticNotebook (Li et al., 2023a) combines data storage/loading with cell replay for optimized session replication. These works do not checkpoint nor restore incrementally, or have limitations/require significant user effort: Dill and ElasticNotebook’s checkpoint files must be entirely loaded for restoration; while KV stores can store and load parts of a state, delta detection and shared reference preservation must be manually handled. In comparison, Kishu can perform low-overhead incremental checkpointing and checkout while preserving shared references (i.e., correctness) and works with almost all data science libraries (§ 7.2).
Memory Snapshotting
@@
There exist OS-level checkpointing tools that can incrementally checkpoint a process for later restoration (CRIU, 2023; Ansel et al., 2009; Chen et al., 1997; Plank et al., 1994; Garg et al., 2018; Jain and Cooperman, 2020; Li and Lan, 2010). These tools identify and store dirty memory pages, then piece together the process image for restoration (CRIU, 2023): they often (1) cause large checkpoint sizes due to coarse page-level deltas (Elnozahy et al., 2002), (2) can only handle single processes (CRIU, 2024), and (3) can only restore from scratch: while we found a patent (Neary et al., 2007) and paper (Ferreira et al., 2011) enabling incremental checkpointing for multiprocessing jobs and with sub-memory-page granularity, respectively, we could not locate working implementations. In comparison, Kishu achieves lower checkpoint overheads via the logical Co-variable granularity deltas (§ 7.3), checkpoint multiprocessing notebooks via application-level instructions (§ 7.2), and incrementally restores via state difference detection and using existing kernel data (§ 7.5).
@@ 9. Conclusion
We have proposed Kishu, a new computational notebook system that offers efficient and fault-tolerant time-traveling between notebook states. Kishu captures session state evolution at a novel Co-variable granularity for efficient incremental checkpointing of state deltas, which Kishu then uses to perform incremental checkout with minimal data loading. Kishu’s contributions include (1) low-overhead state delta detection, (2) branch-based state versioning, and (3) generalizable time-traveling—preserving inter-variable dependencies and handling missing data with fallback recomputation. We have shown that Kishu is compatible with 146 data science object classes and reduces checkpoint storage size and checkout time by up to 4.55$\times$ and 8.18$\times$, respectively, on real-world notebooks.